### v1

# The Comma Sequence is Finite in Other Bases

Abstract
The comma sequence (1, 12, 35, 94, …) is the lexicographically earliest sequence such that the difference of consecutive terms equals the concatenation of the digits on either side of the comma separating them. The behavior of a “generalized comma sequence” depends on the base the numbers are written in, as well as the sequence’s initial values. We provide a computational proof that all comma sequences in bases 3 through 19, as well as 22 and 23, are finite. Relying on a combinatorial conjecture, Angelini et al. estimated that the final element of a comma sequence in base $b$ should be roughly $\exp(O(b))$. We prove their conjecture, but provide evidence that the correct estimate is actually $\exp(O(b\log b))$.
1 Introduction
In 2006, Éric Angelini submitted A121805, the “comma sequence,” to the OEIS [oeis]. It is defined by a peculiar rule, best illustrated by listing the initial terms and their differences:
[…]
The comma sequence is the lexicographically earliest sequence of positive integers such that the difference of consecutive terms equals the concatenation of the digits on either side of the comma separating them. For example, the sequence contains $12,35$, which has a difference of $35-12=23$, which happens to be the concatenation of $2$ (the digit to the left of the comma) and $3$ (the digit to the right of the comma). The amazing fact about the comma sequence is that it contains exactly 2,137,453 terms. It reaches $99999945$ and then terminates.
A (generalized) comma sequence in base $b$ is the lexicographically earliest sequence which satisfies the comma rule with a particular fixed initial value. In [neil], Angelini et al. studied generalized comma sequences and obtained a number of results we will generalize. First, they proved that all comma sequences in base 3 are finite. We have extended this result to more bases.
Theorem 1.
Comma sequences with arbitrary positive initial values are finite in bases 3 through 19, as well as 22 and 23.
The proof of this theorem is a parallelized computation designed in the Go programming language and carried out on a computing cluster. We will give the mathematical details of this proof in Section 3, and the computational details in Section 6. Table 1 provides estimates for the amount of time it would take us to prove that comma sequences in a sampling of bases are finite.
$b$ $L(b)$ estimated time
$3$ 4 461µs
$10$ 924 41ms
$13$ 44324280 5 min
$15$ 1661619960 1 hour
$19$ 70758321480 37 hours
$20$ 493766987313600 6 years
$21$ 373234106765520 5 years
$23$ 430151641920 1 week
Table 1: Estimated runtimes to prove that all comma sequences in select bases $b$ are finite on our local 12-core, 3.7GHz machines. The main runtime consideration is the quantity $L(b)$, defined in (5). Large constant factor improvements can be obtained by further parallelization.
The obvious conjecture is that all comma sequences in all bases are finite. This remains open, but Angelini and company introduced a simple random model which attempts to explain the expected duration of comma sequences in different bases. They made a combinatorial conjecture about this model, which we have confirmed.
Theorem 2.
The number $D(b)$ of comma sequences in base $b\geq 2$ with initial values in $[b^{m}-b^{2},b^{m})$ (with $m\geq 2$) that do not reach $b^{m}$ is independent of $m$ and has ordinary generating function
$\sum_{b=2}^{\infty}D(b)t^{b}=(1-t)^{-1}\left(\sum_{b=1}^{\infty}\frac{t^{b(b+3)/2}}{(1-t^{b})}-t^{2}\right)=t^{3}+2t^{4}+4t^{5}+5t^{6}+\cdots$ (1)
On the basis of this theorem, the authors of [neil] (with help from Václav Kotĕs̆ovec) argued that the final element of a base $b$ comma sequence should be roughly $e^{2b}$. More precisely, they estimated that a comma sequence ought to survive about $2b/\log(2b)$ “danger intervals,” which we will define below. In Section 5, we will define a different probabilistic model which predicts that comma sequences should survive about $b/2+1$ intervals, which translates to an estimate length of $b^{b/2+1}=e^{O(b\log b)}$. This estimate is more accurate for large bases, but depends on a probabilistic conjecture that we cannot prove. We hope that some expert can translate at least one of these models into a rigorous argument.
2 Background and examples
Because comma sequences and their generalizations are not widely known, in this section we summarize some relevant background [neil] and give a number of examples. We begin with a more explicit definition of generalized comma sequences.
Definition 1.
The generalized comma sequence in base $b$ with initial condition $v$ is the sequence $a(n)$ defined as follows: $a(1)=v$; for $n>1$, if $x$ is the least significant base $b$ digit of $a(n-1)$, then
$a(n)=a(n-1)+bx+y,$ (2)
where $y$ is the most significant digit of $a(n)$ and is the smallest such $y$, if any exists. If no such $y$ exists, then the sequence terminates.
The comma sequence corresponds to $v=1$ and $b=10$. The first ten terms of the base $10$ comma sequences with initial values $v=1,2,\dots,10$ are as follows:
$1,12,35,94,135,186,248,331,344,387\dots$
$2,24,71,89,180,181,192,214,256,319\dots$
[…]
$4,48,129,221,233,265,318,402,426,490\dots$
$5,61,78,159,251,263,295,348,432,456\dots$
$6,73,104,145,196,258,341,354,397,471\dots$
$7,85,136,197,269,362,385,439,534,579\dots$
$8,97,168,250,252,274,317,390,393,427\dots$
$9,100,101,112,133,164,206,268,351,364\dots$
$10,11,23,58,139,231,243,275,328,412\dots$
These sequences are all finite. Their lengths are recorded in A330128.
The first ten terms of the base $3$ comma sequences with initial values $v=1,2,3$ are as follows (given in base $3$):
$(1)_{3},(12)_{3},(110)_{3},(111)_{3},(122)_{3},(221)_{3},(1002)_{3},(1100)_{3},(1101)_{3},(1112)_{3}\dots$
[…]
These sequences are also all finite; the first lasts for seventeen terms.
The numbers which do not have successors, meaning integers for which comma sequences terminate, are called “landmines.” They are precisely the integers in base $b$ of the form
$(b-1,b-1,\dots,b-1,x,y)_{b}$
where $x$ and $y$ are nonzero digits which sum to $b-1$ [neil, Theorem 5.2]. (We use the notation $(d_{m},d_{m-1},\dots,d_{0})_{b}$ to represent the integer $d_{m}b^{m}+d_{m-1}b^{m-1}+\cdots+d_{0}$ in base $b$ where $d_{i}$ are digits base $b$.) For example, in base $10$, the integers 45, 99972, and 9999918 are all landmines.
Landmines live in the following small intervals just to the left of powers of $b$.
Definition 2.
The base $b$ danger intervals are the intervals $[b^{m}-b^{2},b^{m})$ for $m\geq 2$.
Given a positive integer $x$, call the set of integers which could follow $x$ in a base $b$ comma sequence $F_{b}(x)$. A landmine is characterized by $|F_{b}(x)|=0$. In general, $|F_{b}(x)|\leq 2$. To see this, note that the next term of a comma sequence following $x$ is $x+d$ for some two digit number $d$, where the most-significant digit of $d$ is determined by $x$. That leaves only two possibilities: a carry occurs or does not. In our “strict” definition, we resolve the possible ambiguity of which $d$ to pick by requiring that the smallest possible successor is chosen.
In the other direction, it can be shown that that every integer greater than $b^{2}$ belongs to precisely one $F_{b}(x)$. Of them, the only ones which are never chosen as a successor in a base $b$ comma sequence are of the form
$(d,0,0,\dots,0)_{b},$
where $d>1$ is a base $b$ digit [neil, Theorems 5.4, 5.5].
From a more graphical perspective, imagine a directed graph $C_{b}$ on the positive integers which contains the edge $i\to j$ provided that $j$ would satisfy (2) (though would perhaps be the larger of two options) in base $b$. What we have just said amounts to the observation that $C_{b}$ has out-degree at most $2$ and in-degree at most $1$, and further that $C_{b}$ is a union of finitely many disjoint (directed) trees which partition the integers. As a corollary, it contains an infinite path, meaning that if comma sequences were allowed to pick either of two possible successors, then there would be at least one infinite comma sequence in every base [neil, Theorem 5.6].
3 The computational proof
Neil Sloane and Giovanni Resta communicated to us the rough idea behind their proof that all comma sequences in base $3$ are finite. We will describe their idea and its generalization to arbitrary bases, illustrate some optimizations that we made, and give conjectured runtimes for applying the algorithm in arbitrary bases. This section contains primarily mathematical details, while Section 6 contains more details about the computation and its implementation.
Definition 3.
Given a base $b$ digit $d$, an integer $0\leq u<b^{2}$, and a positive integer $k$, we define $(d,u,k)$ to represent the integer $d\cdot b^{k}-u$.
Comma sequences can be thought of as the result of repeatedly applying a certain complicated map to the points $(d,u,k)$. For example, in the original comma sequence, we see the values $a(4)=94$, $a(5)=135$, later $a(6)=186$, and then $a(7)=248$. We would describe this as
$(1,6,2)\to(2,14,2)\to(3,52,2).$
Note that we omitted $a(5)=135$ in this diagram. This is because we always take the smallest possible value of $u$ for a given $d$ and $k$, and any comma sequence in base 10 which reaches $135$ (alias $(2,65,2)$) will then reach $186$ (alias $(2,14,2)$). More generally, for any base $b$ and digit $d$, there is a finite set of “minimal $u$’s” which all other $u$’s lead to, and these values are independent of $k$. The precise statement of this fact and its (annoyingly technical) proof are the following lemma.
Lemma 1 (Minimal $u$’s).
For any base $b>2$, any integer $k>2$ and any nonzero base $b$ digit $d$, there exists a smallest finite set $U(b,d)$ that satisfies the following property: If any base $b$ comma sequence has a value in $[d\cdot b^{k}-b^{2},d\cdot b^{k})$, then its final value in this interval is $d\cdot b^{k}-u$ for some $u\in U(b,d)$.
If $d\neq 1$, then
[…]
If $d=1$, then
[…]
Proof.
Begin with $d>1$. The final value of a base $b$ comma sequence in $[d\cdot b^{k}-b^{2},d\cdot b^{k})$ will have the form $(d-1,b-1,\dots,b-1,x,y)_{b}$ for some $0\leq x,y<b$. Because this is the last term in the interval, the next term is
$(d-1,b-1,\dots,b-1,x,y)+(y,d)_{b},$
The relationship between $(x,y)$ and $(r,s)$ is
$(x,y)_{b}+(r,s)_{b}=b^{2}.$
There are two cases: First, $x+y\geq b$. Since $y>0$, we have $(r,s)=(b-1-x,b-y)$ and the inequality $x+y\geq b$ is equivalent to $r+s\leq b-1$ and $s>0$.
Second, $x+y=b-1$ and $y+d\geq b$. Note that
$(d-1,b-1,\dots,b-1,x,y)+(y,d-1)_{b}$
is also no longer in the interval, so $y+d-1\geq b$. This again implies $y>0$, and so $(r,s)=(b-1-x,b-y)$, and the conditions
$x+y$ $=b-1$
$y+d$ $\geq b+1$
are equivalent to
$r+s$ $=b$
$s$ $<d.$
The reasoning for $d=1$ is similar. The only difference is in the second case, when the inequality $y+d-1\geq b$ becomes $y+b-1\geq b$. The rest of the argument is identical. ∎
Definition 4.
The digraph $G_{b}$ consists of vertices labeled $(d,u,k)$ with base $b$ digit $d$, a minimal $u$ with respect to $d$, and an integer $k\geq 0$. The edge $(d,u,k)\to(d^{\prime},u^{\prime},k^{\prime})$ exists if the latter is the immediate image of the former under the comma map.
Our main goal is to prove that $G_{b}$ does not have infinite paths. In the remainder of this section we will show how this reduces to a finite problem. To begin, we must describe a computational shortcut.
Computing the transform
Given $(d,u,k)$ which is not a landmine, we want to determine the relation
$(d,u,k)\to(d^{\prime},u^{\prime},k^{\prime}).$
For example, if we have a base $10$ sequence at the value $(6,8,3)$ (also known as $5992$), where does it go next? Both $d^{\prime}$ and $k^{\prime}$ are easy: $d^{\prime}=d+1$ if $d<b-1$, and $d^{\prime}=1$ otherwise. Similarly, $k^{\prime}=k$ if $d<b-1$, and $k^{\prime}=k+1$ otherwise. So, as a partial answer, $(6,8,3)\to(7,u^{\prime},3)$. The main difficulty is determining $u^{\prime}$.
Starting from $5992$, the next few differences of the sequence would be $26,86,46,06$, and $66$. After that, they cycle through the same values until the comma sequence reaches a number which begins with a $6$. In general, if our current value is $(d,u,k)$, then Angelini and friends proved that the difference between consecutive terms cycle through the arithmetic progression
$S_{b}(d,u)=\{((md-u)\bmod b)b+d\mid 0\leq m<b/\gcd(b,d)\}$ (3)
until they reach $(d^{\prime},u^{\prime},k^{\prime})$ (see [neil, sec. 6] and A121805). Therefore, to compute $u^{\prime}$, we start at $d\cdot b^{k}-u$ and add the elements of $S_{b}$ in order, repeating until we get as close to $(d+1)b^{k}$ as possible without going over.
The distance between $(d+1)b^{k}$ and $db^{k}-u$ is $b^{k}+u$. Using entire copies of $S_{b}(d,u)$, we can get as close as
$(b^{k}+u)\bmod\sum S_{b}(d,u)$ (4)
where $\sum S_{b}(d,u)$ is the sum of the entries of the cycle. After this, we must manually add elements of $S_{b}(d,u)$, starting from the beginning, to close the gap.
Making the problem finite
The key point is (4). It is well-known that $b^{k}\bmod m$, for integers $b$ and $m\geq 1$, is eventually periodic in $k$. (Indeed, any C-finite sequence taken with a fixed modulus is eventually periodic [pisano, carmichael, concrete].) This means that, for each cycle $S_{b}(d,u)$, there exists a period $l(d,u)$ and a nonnegative integer $k_{0}(d,u)$ such that
$b^{k}\equiv b^{k+l(d,u)}\bmod{\sum S_{b}(d,u)}$
for all $k\geq k_{0}(d,u)$. In particular, for sufficiently large $k$ we only need to compute (4) for values of $k\bmod l(d,u)$. If we set
$L(b)=\mathrm{lcm}\{l(d,u)\mid 1\leq d<b,\quad u\in U(b,d)\cup\{0\}\},$ (5)
then $b^{k}\bmod\sum S_{b}(d,u)$ is eventually periodic with period $L(b)$ for all $(d,u)$. This implies that, to determine the behavior of comma sequences with sufficiently large $k$, we only need to consider points $(d,u,k)$ where $k$ is the representative of an equivalence class modulo $L(b)$.
The technical point of forcing $k$ to be “sufficiently large” is not important for our proof. If there were a truly infinite comma sequence, then its values would pass by arbitrarily large $b^{k}$’s, meaning that the periodicity of (4) would kick in at some point.
Definition 5.
The digraph $G_{b}^{\prime}$ consists of vertices $(d,u,\kappa)$ where $d$ is a base-$b$ digit, $u$ is an element of $U(b,d)$, and $[\kappa]$ is an equivalence class mod $L(b)$. The edge $(d,u,[\kappa])\to(d^{\prime},u^{\prime},[\kappa^{\prime}])$ exists provided that the edge $(d,u,k)\to(d^{\prime},u^{\prime},k)$ exists in $G_{b}$ for sufficiently large integers $k\equiv\kappa\pmod{L(b)}$ and $k^{\prime}\equiv\kappa^{\prime}\pmod{L(b)}$.
Lemma 2.
$G_{b}^{\prime}$ is well-defined and has both in- and out-degree at most $1$.
Proof.
The periodicity of (4) for sufficiently large $k$ shows that the computation of $(d,u,k)\to(d^{\prime},u^{\prime},k^{\prime})$ in $G_{b}$ depends only on $k\bmod L(b)$, so $G_{b}^{\prime}$ is well-defined. It further shows that the in- and out-degree of $(d,u,[\kappa])$ in $G_{b}^{\prime}$ will be the same as $(d,u,k)$ in $G_{b}$ for sufficiently large $k\equiv\kappa\pmod{L(b)}$. Vertices in $G_{b}$ have out-degree at most $1$ (by the lexicographically earliest construct) and in-degree at most $1$ (by Lemma 5.3 in [neil]). ∎
Proposition 1.
Comma sequences with arbitrary initial conditions in base $b$ terminate if and only if $G_{b}^{\prime}$ contains no cycles.
Proof.
The existence of infinite comma sequences is equivalent to the existence of infinite walks on the vertices of $G_{b}$, which is in turn equivalent to the existence of a cycle in $G_{b}^{\prime}$. ∎
Lemma 2 implies that $G_{b}^{\prime}$ is the disjoint union of independent paths and cycles. Theorem 5.4 of [neil] says that the only points with in-degree 0 are $(d,0,k)$ with $d>1$, so these must be the start of the disjoint paths in $G_{b}^{\prime}$. Thus, to check that $G_{b}^{\prime}$ contains no cycles, we can iterate the comma map on the points $(d,0,k)$ to generate all paths, then check that we are not missing any vertices. We carried out this computation for $b\in\{3,4,\dots,19\}\cup\{22,23\}$ and found that $G_{b}^{\prime}$ consisted of only paths, thereby proving that all comma sequences in these bases are finite.
4 The geometric model
Definition 6.
Let $D(b)$ be the number of comma sequences in base $b\geq 2$ with initial values in $[b^{k}-b^{2},b^{k})$ (with $k>2$) that do not reach $b^{k}$. (As mentioned in the previous section, this quantity is independent of $k$.)
The probabilistic model of [neil] makes the simplifying assumption that comma sequences enter the danger intervals $[b^{k}-b^{2},b^{k})$ uniformly randomly, and that each interval should be considered independent. Under this model we would expect to hit a landmine after $b^{2}/D(b)$ intervals. In this section we will prove that $D(b+1)-D(b)$ is essentially A136107, the number of ways to write $b+1$ as the difference of positive triangular numbers. This will establish the generating function identity (1), and as a corollary the asymptotic estimate
$\frac{b^{2}}{D(b)}\sim\frac{2b}{\log 2b}.$
While this result is correct, the underlying model is slightly off. See Section 5 for better estimates.
For brevity, we will refer to the base $b$ integer $(b-1,b-1,\dots,b-1,x,y)_{b}$ as $\{x,y\}_{b}$.
Lemma 3.
For $y>0$, $\{x,y\}_{b-1}$ is the parent of $\{r,s\}_{b-1}$ in a comma sequence for base $b-1$ if and only if $\{x+1,y\}_{b}$ is the parent of $\{r+1,s\}_{b}$ in base $b$.
Proof.
By the comma rule, we have
$\{r,s\}_{b-1}=\{x,y\}_{b-1}+(y,b-2)_{b-1}.$
There must be a carry in the units digit because $y$ is positive and $s$ is less than $b-1$, so $y+(b-2)=s+b-1$. Because there is no carry in the tens digit this gives $x+y+1=r$. Together these imply $y+(b-1)=s+b$ and $(x+1)+y+1=(r+1)$, so $\{x+1,y\}_{b}$ is followed by $\{r+1,s\}_{b}$ in base $b$. (Note that $r+1$ and $x+1$ are digits base $b$ since $r$ and $x$ are digits in base $b-1$.) The other direction follows the same logic. ∎
Lemma 4.
$D(b)-D(b-1)$ is one more than the number of elements of the form $\{0,d\}_{b}$ with $10<d<b-$ 1 which fail to escape the base $b$ interval.
Proof.
If $\{r,s\}_{b-1}$ is a landmine in base $b-1$, then $\{r+1,s\}_{b}$ is a landmine in base $b$. By Lemma 3, if $x,y>0$ and $\{x,y\}_{b-1}$ leads to the landmine $\{r,s\}_{b-1}\neq\{x,y\}_{b-1}$ in base $b-1$, then $\{x+1,y\}_{b}$ leads to $\{r+1,s\}_{b}$ in base $b$. Together these account for $D(b-1)$ failed escapes. Elements of the form $\{x,0\}_{b}$ are followed by $\{x,b-1\}_{b}$ (with a difference of $(0,b-1)_{b}$) and those escape with a difference of $(b-1,1)_{b}$. Therefore, the only possible “new” escapes are (1) the landmine $\{1,b-2\}_{b}$ not covered by the map $\{r,s\}_{b-1}\mapsto\{r+1,s\}_{b}$ and (2) elements of the form $\{0,y\}_{b}$ with $10<y<b-$ 1. ∎
Lemma 5.
With $10<y<b-$ 1, the element $\{0,y\}_{b}$ fails to escape the interval in base $b$ if and only if there is a positive integer $n$ such that
$b=\frac{(n+1)(2(y+1)-n)}{2},$ (6)
with a single exception if $b$ is a triangular number. If $b=(k+1)(k+2)/2$ for some integer $k$, then the element $\{0,k\}_{b}$ escapes despite the solution $n=y=k$.
Proof.
If we continue the comma sequence beginning at $\{0,y\}_{b}$, then as long as we are in the interval we obtain the sequence of values $\{x(n),y(n))\}_{b}$ which satisfy $\{x(0),y(0)\}_{b}=\{0,y\}_{b}$ and the recurrences
$y(n+1)$ $=y(n)-1$
$x(n+1)$ $=x(n)+y(n)+1.$
From these recurrences it is easy to prove
$x(n)+y(n)=(n+1)y-\frac{n(n-1)}{2},$
and so $x(n)+y(n)=b-1$ is equivalent to (6).
If $\{0,y\}_{b}$ hits a mine in $n$ steps then $x(n)+y(n)=b-1$, which proves half of our lemma. For the converse, suppose that there is a positive integer $n$ for which (6) has a solution. Then we have $x(n)+y(n)=b-1$, but this does not mean that the sequence terminates. The possibilities are as follows:
  1. 1.
    We hit a mine.
  2. 2.
    We hit $\{0,b-1\}_{b}$. We excluded this with $d<b-1$.
  3. 3.
    We hit $\{b-1,0\}_{b}$. In this case, we have $y(n)=n-y=0$, so $n=y$. Plugging this into (6) shows that
    $b=\frac{(n+1)(n+2)}{2}.$
    Therefore when $b$ is a triangular number, there is precisely one $y$ which leads to a solution to (6), but should not be counted.
It follows that solutions to (6) determine all elements $\{0,y\}_{b}$ which fail to escape the interval, except for when $b$ is triangular. ∎
Lemma 6.
The number of $10<y<b-$ 1 for which there are positive integer solutions to (6) is one less than the number of odd divisors of $b$.
Proof.
Note that (6) is equivalent to the following quadratic in $n$:
$n^{2}-(2y+1)n+2(b-1-y).$
Because the linear term is odd and the constant term is positive, for any $y$ for which there are integer solutions, there will in fact be two, one even and one odd. Thus the number of $y$ which have solutions is in correspondence with the number of even $n$ which lead to integer solutions. If $n$ is even then (6) holds if and only if $n+1$ divides $b$, which establishes a correspondence with the odd divisors of $b$. Note that $b-1$ will always be a solution and 00 will never be. ∎
Proposition 2.
$D(b)-D(b-1)=A136107(b)$ for $b\geq 3$.
Proof.
By Lemma 4, $D(b)-D(b-1)$ is one more than the number of elements $(0,d)$ which fail to escape the interval in base $b$. Lemmas 5 and 6 tells us that this is the number of odd divisors of $b$, except when $b=k(k-1)/2$ is a triangular number, in which case we must discard a single divisor. According to a comment by R. J. Mathar this is equivalent to $A136107(b)$. ∎
Proof of Theorem 2.
The previous proposition implies
$D(b)=\sum_{k=1}^{b}A136107(b)-1$
for $b\geq 2$. According to a comment by Vladeta Jovovic in the OEIS, the ordinary generating function for A136107 is
$\sum_{n\geq 1}\frac{t^{n(n+3)/2}}{1-t^{n}}.$
Using the formula for the generating function of partial sums [wilf], we see that the ordinary generating function of $D(b)$ is
$\frac{1}{1-t}\sum_{n\geq 1}\frac{t^{n(n+3)/2}}{(1-t^{n})}-\frac{t^{2}}{1-t},$
which matches (1). ∎
5 A more accurate model
In Section 3 we constructed a finite graph $G_{b}^{\prime}$ which contains no cycles if and only if all comma sequences in base $b$ are finite. By randomly sampling points from this graph, we obtain a different probabilistic model of comma sequence behavior. This model predicts the lengths of comma sequences more accurately than the geometric model in [neil], provided that there truly are no cycles.
The expected number of danger intervals which a comma sequence survives can reasonably be computed as
$|V(G_{b}^{\prime})|^{-1}\sum_{v\in V(G_{b}^{\prime})}h(v),$
where $h(v)$ is the number of edges between $v\in V(G_{b}^{\prime})$ and the end of the path in which it resides, and $V(G_{b}^{\prime})$ is the set of all vertices in $G_{b}^{\prime}$. A simple calculation shows that this equals
$|V(G_{b}^{\prime})|^{-1}\sum_{P}{|P|\choose 2},$
where the sum is over all paths $P$ in $G_{b}^{\prime}$ and $|P|$ is number of vertices in a path. By adjusting the denominator, we can rephrase this as an expectation in terms of a uniformly randomly chosen path $P$:
[…]
Recall that all paths in $G_{b}^{\prime}$ begin at a point of the form $(d,0,\kappa)$ with $\kappa$ taken mod $L(b)$ (as in (5)), so there are $(b-2)L(b)$ paths in total.
Empirical testing led us to the following conjecture.
Conjecture 1.
As $b\to\infty$, the path length $|P|$ is approximately exponentially distributed with rate close to $(b/2+1)^{-1}$. In particular, $\mathbb{E}[|P|]\approx b/2+1$ and $Var(|P|)\approx(b/2+1)^{2}$.
With this conjecture and the preceding remarks, we may very roughly estimate the number of danger intervals survived as
$\frac{b}{2}+1+O(b^{-1}).$ (7)
The geometric model of [neil] predicts that roughly $2b/\log(2b)$ intervals would be survived. Figure 1 gives evidence that $b/2+1$ is a better estimate. It also suggests that $b/2+1$ may also be slightly smaller than the true mean.
Figure 1: Average number of danger intervals survived for comma sequences in base $b$, estimated empirically using the initial values $1$, $2$, …, $b^{2}-1$. Note that the empirical curve begins close to $2b/\log(2b)$, but pulls away for later bases.
6 Computational details
The considerations of Section 3 lead to Algorithm 1. Note that the runtime depends directly on $L(b)$. This quantity begins small ($L(3)=4$ allowed $b=3$ to be done by hand in [neil]) but grows very rapidly. In this section we will mention some practical details of the implementation necessary to handle such large graphs.
Algorithm 1 Check whether base $b$ comma sequences are finite.
procedure IsFinite($b$)
     $L\leftarrow L(b)$
     count $\leftarrow 0$
     for $2\leq d<b$, $0\leq k<L$ do
         $P\leftarrow(d,0,k)$
         repeat
              if $d=0$ then
                  count $\leftarrow$ count$+1$
              end if
              $P\leftarrow$ AdvancePoint($b$, $P$)
              $(\ast,u,\ast)\leftarrow P$
         until isMine($u$)
     end for
     return count == ValidU($b$) $\times$ L
end procedure
procedure ValidU($b$) $\triangleright$ counts valid $(1,u)$ pairs
     count $\leftarrow 0$
     for $1\leq u<b^{2}$ do
         if $u\leq b(-u\bmod b)+b-1$ then
              count $\leftarrow\text{count}+1$
         end if
     end for
     return count
end procedure
procedure AdvancePoint($b,d,u,k$)
     $S\leftarrow\textsc{Cycle}(d,u)$triangleright see (3)
     $(k_{0},l)\leftarrow\textsc{Order}(b,\sum S)$triangleright $b^{k}$ is periodic with period $l$ for $k\geq k_{0}$
     $gap\leftarrow(b^{k_{0}+(k-k_{0})\bmod l}+u)\bmod\sum S$
     if $gap=0$ then
         $gap\leftarrow\textsc{LastElem}(S)$triangleright the shortcut does not apply if the gap is 0
     else
         $i\leftarrow 0$
         while $gap>S[i]$ do
              if $gap<b^{2}$ and $d=b-1$ and IsMine($b$, $gap$then
                  break
              end if
              $gap\leftarrow gap-S[i]$
              $i\leftarrow i+1$
         end while
     end if
     $u^{\prime}\leftarrow gap$
     if $d=b-1$ then
         return $(1,u^{\prime},(k+1)\bmod L(b))$
     else
         return $(d+1,u^{\prime},k)$
     end if
end procedure
Runtime and memory
The algorithm hopes to see every point of the form $(d,u,k)$ where $d>0$ is a base $b$ digit, $u$ is in $U(b,d)$, and $k$ is taken mod $L(b)$. The set $U(b,d)$ has size roughly (${b\choose 2}$, so the number of points to process is roughly $b^{3}L(b)/2$. There are a number of improvements to be made.
First, the paths of $G_{b}^{\prime}$ are processed independently, so the procedure IsFinite can be trivially parallelized. Second, the procedure AdvancePoint duplicates effort over repeated calls. The computations in that procedure depends only on $d$, $u$, and $k\bmod l(d,u)$, the periods appearing in Equation (5). So, while IsFinite must consider $k\bmod L(b)$, we can cache the results of AdvancePoint(d, u, k mod l(d, u)) and save a lot of time. This cache requires a memory footprint of roughly
$O\left((\log b)^{3}\sum_{d,u}l(d,u)\log l(d,u)\right),$
which is large but has not been a bottleneck thus far. (For $b=23$, we used 300 parallel processes which each stored their own cache. The total memory consumption was approximately 34GB.)
We implemented a parallel version of the algorithm given here in the Go programming language [^1] As a warning, this design seems particularly poorly suited to Python, thanks to its Global Interpreter Lock [gil].[go, github]. The practical details of the parallelism are not entirely trivial; see Figure 2.
Figure 2: Diagram of parallel Go implementation of Algorithm 1. Solid lines indicate (possibly buffered) channel communication. Dashed lines indicate non-blocking memory access or atomic operations. On large compute clusters, it may be desirable to use several copies of this program with independent caches to reduce the overhead of inter-processor communication.
As a minor remark, it proved difficult to find a good reference algorithm for determining precisely when and with what period $b^{k}\bmod m$ repeats for arbitrary integers $b$ and $m\geq 1$. For completeness, this is how it is done:
  • If $\gcd(b,m)=1$, then the period equals the smallest positive $k$ such that $b^{k}\equiv 1\pmod{m}$, and the cycle begins immediately [hardywright, ch. 6].
  • If $\gcd(b,m)=g>1$, then let $k_{0}$ be the smallest positive integer such that $\gcd(b^{k},m)$ is constant for $k\geq k_{0}$. Then $b^{k}\equiv b^{k+l}\pmod{m}$ for $k\geq k_{0}$ if and only if $b^{l}\equiv 1\pmod{m/g}$. Because $b$ and $m/g$ are relatively prime, the smallest such $l$ can be determined by the previous argument. To see that this gives the smallest such $k_{0}$, note that if $b^{k}\equiv b^{k+l}\pmod{m}$ for all $k\geq k_{1}$, then $\gcd(b^{k},m)=\gcd(b^{k+l},m)$ shows that $\gcd(b^{k},m)$ is constant for $k\geq k_{1}$. Thus $k_{1}\geq k_{0}$.
Limitations
In practice, the runtime of our program seems to follow a fairly exact power law. The number of seconds required to process base $b$ on our local machines is on the order of $L(b)^{1.01}$. The constant factor can be dramatically improved by parallelization, but $L(b)$ grows fast enough that constant factor improvements will probably not be enough to cover more bases than we have done here using our method. Table 1 (page 1) gives some explicit runtime estimates for larger bases.
Acknowledgements
We thank Neil Sloane and Giovani Resta for bringing this problem and several key ideas to our attention. We also acknowledge the Office of Advanced Research Computing (OARC) at Rutgers, The State University of New Jersey for providing access to the Amarel cluster and associated research computing resources that have contributed to the results reported here. URL: https://oarc.rutgers.edu.
[…]